Quiz: Quiz - Chapter 3 7 of 12 (3 This Quiz: 12 pts 0 257 587 0 (a) Check the data set for outiers 389 3027 59 (c) Provide an explanation for any outlers 71 184 332 450 A. O B. The outiler(s) is are □ There are no outiers (Use a comma to separate answers as needed ) OA. 5 6 8
Answer #1

Here for detecting outlier we will first write the table in ascending order

0
0
0
65
71
151
184
226
257
280
304
320
332
355
389
459
567
587
595
3027

Here as n = 20

Qquartile 1 = Q1 = (5th number + 6th number)/2 = (71 + 151)/2 =  111

Quartile 2 = Q3 = (15th number + 16th number) = (389+459)/2 = 424

Inter quartile range = 424 -111 = 313

so here as we know that any number will be an outlier if it is outside the limits of Q1 - 1.5 IQR or Q3 + 1.5 IQR

so here

Q1 - 1.5 IQR = 111 - 1.5 * 313 = -358.5

Q3 + 1.5IQR = 424 + 1.5 * 313 = 893.5

so here as we see that number 3027 is an outlier as it is outside the limit.

(b) Here option b graph is correct.

(c) Here part (c) is not given.